OpenStudy (anonymous):

I WILL GIVE METAL TO THE FIRST 2 PEOPLE WHO HELPS ;) - x^2-22x+121& x^2+24x+144

OpenStudy (anonymous):

i'll give you a medal if you write what the question is at the moment you have written two polynomials without a question

OpenStudy (anonymous):

it says to find the factor and polynomials

OpenStudy (anonymous):

\[x^2+24x+144\] is a perfect square, since \(144=12^2\) and \(24=2\times 12\) so \[x^2+24x+144=(x+12)^2\]

OpenStudy (anonymous):

\[ -x^2-22x+121\] does not factor using whole numbers, although\[ x^2-22x+121\]does

OpenStudy (anonymous):

just like the other one since \(121=11^2\) and \(-22=-2\times 11\) you have \[x^2-22x+121=(x-11)^2\]